OpenStudy (anonymous):

2. a. graph y = -2 | x - 1 | + 4

OpenStudy (anonymous):

since the equation is a linear equation, y = mx+b, we can do this: find x-intercept by substituting 0 to y and equate for x. find y-intercept by substituting 0 to x and equate for y. x-intercept: 0 = -2|x-1|+4 x = 1 so point 1 is (1,0) y-intercept: y = -2|0-1|+4 y = 2 so point 2 is (0,2) now plot the points..

OpenStudy (anonymous):

thanks you were a lot of help :)

OpenStudy (anonymous):

no problem. just practice, it helps a lot too :)